# Estimating Semantic Alphabet Size for LLM Uncertainty Quantification

### v1

@@ Appendix D Additional incorrectness detection results
Detailed results for incorrectness classification across three datasets, four LLMs, and ten UQ methods are shown in Figure 7. Error bars are $95\%$ CIs about empirical AUROC values, calculated via DeLong’s method (DeLong et al., 1988; Sun & Xu, 2014).
@@
Figure 7: Performance of incorrectness classification on single-answer QA for uncertainty measures using $n=10$ samples. “Plugin,” “CS-GT,” and “CS-H” are estimators of semantic entropy. “NumSets,” $U_{EigV}$,” and $\widehat{|S|}_{Hybrid}$” are estimators of semantic alphabet size. Error bars reflect $95\%$ CIs about the AUROC.
Appendix E Consistency in evaluation
@@
The relevant LLM UQ literature is not always consistent in experimental setup. For instance, the plurality of considered works use $n=10$ sampled responses per query (Kuhn et al., 2023; Farquhar et al., 2024; Nikitin et al., 2024; Nguyen et al., 2025), but Lin et al. (2024) use $n=20$. Some works vary prompting templates across datasets (e.g., zero-shot vs. multi-shot prompting) (Kuhn et al., 2023; Lin et al., 2024). A variety of methods for automated LLM incorrectness evaluation have been employed, including thresholded ROUGE-L (Kuhn et al., 2023), binary LLM-as-judge (Farquhar et al., 2024), thresholded consistency-based LLM-as-judge Lin et al. (2024), and BERTScore (Nguyen et al., 2025). Although it is out of this work’s scope to exhaustively ablate all such choices, we attempt to be explicit in describing our selections to support future work.

### v2

@@ Appendix F Additional incorrectness detection results
Detailed results for incorrectness classification across three datasets, five LLMs, and eleven UQ methods are shown in Figure 10. Error bars are $95\%$ CIs about empirical AUROC values, calculated via DeLong’s method (DeLong et al., 1988; Sun & Xu, 2014).
@@
Figure 10: Performance of incorrectness classification on single-answer QA for uncertainty measures using $n=10$ samples. Methods considered include black-box semantic alphabet size estimators ($\widehat{|S|}_{Hybrid}$, NumSets, $\widehat{|S|}_{GT}$, $U_{EigV}$), explicit discrete semantic entropy (DSE) estimators ($\hat{\mathbb{H}}_{Hybrid},\hat{\mathbb{H}}_{Plugin}$, $\hat{\mathbb{H}}_{CS}$), and other uncertainty estimators, which include white-box (PE, SE) and black-box methods (SNNE, KLE). Error bars reflect $95\%$ CIs about AUROC, via DeLong’s method.
Appendix G Underestimation of semantic alphabet size
In Figure 2 we exhibit underestimation of SE, using the ratio of DSE estimators (using varying sample sizes) to $SE^{*}$. We provide a similar illustration in Figure 11: NumSets, the implicit alphabet size estimator of canonical DSE, underestimates the number of semantic categories for typical sample sizes, and the the hybrid semantic alphabet size estimator does so by less.
Figure 11: Illustrating underestimation in alphabet size estimation with typical sample sizes. The ratios of alphabet size estimators with varying sample size ($n=5,10,25,50,75,100$) to the observed number of semantic categories with $n=100$ (denoted $S^{*}$) are shown, with values below 11 suggesting underestimation (dotted grey line). The estimators displayed are NumSets, used by the plugin estimator of canonical discrete semantic entropy (dotted indigo line) and the “hybrid” alphabet size estimator of Equation 9 (solid indigo line). Results are averaged across queries within each dataset, then uniformly averaged across datasets. Log scale is used on the x-axis to highlight differences between estimators with smaller sample sizes.
Appendix H Implementation variations
@@
The relevant LLM UQ literature is not uniform in experimental setup. For instance, the plurality of considered works use $n=10$ sampled responses per query (Kuhn et al., 2023; Farquhar et al., 2024; Nikitin et al., 2024; Nguyen et al., 2025), but Lin et al. (2024) use $n=20$. Some works vary prompting templates across datasets (e.g., zero-shot vs. multi-shot prompting) (Kuhn et al., 2023; Lin et al., 2024). A variety of methods for automated LLM incorrectness evaluation have been employed, including thresholded ROUGE-L (Kuhn et al., 2023), binary LLM-as-judge (Farquhar et al., 2024), thresholded consistency-based LLM-as-judge Lin et al. (2024), and BERTScore (Nguyen et al., 2025). Although it is out of this work’s scope to exhaustively ablate all such choices, we attempt to be explicit in describing our selections to support future work.